OpenStudy (zenmo):

Chain rule: if F(x) = f(xf(xf(x))), where f(1) = 2, f(2) = 3, f ' (1) = 4, f ' (2) = 5, and f ' (3) = 6, find F ' (1).

OpenStudy (owen3):

Is this the question$$F(x) = f(f(f(x)))$$

OpenStudy (owen3):

\[F(x) = f(f(f(x)))\]$$ F'(x) = f'(f(f(x)))\cdot f'(f(x)) \cdot f'(x) $$

OpenStudy (zenmo):

the book has the question as stated

OpenStudy (zenmo):

but i think yours is the same still

OpenStudy (zenmo):

is the answer 120?

OpenStudy (owen3):

The notation matters \[ f(xf(xf(x))) \neq f(f(f(x)))\]

OpenStudy (zenmo):

\[F'(1) = f'(f(2)) \times f'(2) \times 4 = 120\]

OpenStudy (owen3):

yes that's correct, but the x's in your question are not superfluous. I think it's meant to be read differently

zepdrix (zepdrix):

Ya this is a huge product rule mess if it's stated correctly :) lol